# Maximal work extraction unitarily from an unknown quantum state: Ergotropy estimation via feedback experiments

### v1

@@ I Introduction
Quantum technologies are appealing as they can surpass their classical counterparts by using exclusive concepts like coherence and entanglement. Understanding their capabilities has progressed in recent years, thanks to the numerous experimental breakthroughs that have improved control over quantum states. Specifically, the investigation of the principles of quantum thermodynamics is made possible by quantum thermal machines, which include refrigerators and heat engines to regulate heat flow and work production [1, 2]. Further options for storing energy for later extraction include quantum batteries [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. In every possible framework, like charging speed [15, 16, 17, 18, 19, 20], stored energy [21, 22], and energy extraction [23, 24, 25, 26, 27], the quantum batteries have the potential to outperform their classical counterparts.
@@
The main focus of this work is FQErgo, an algorithm we propose for preparing the passive state and thereby estimating ergotropy. FQErgo is inspired by (hence the name) the feedback-based algorithm for quantum optimization (FALQON) [42], which has been recently developed for combinatorial quantum optimization. This method uses feedback conditioned on qubit measurements at each quantum circuit layer and parametrized quantum circuits to determine the values of the circuit parameters at subsequent layers. FALQON is related to quantum Lyapunov control (QLC), a feedback procedure to find the controls needed to manipulate a quantum system’s dynamics in a desired way [43, 44].
We note several salient features of the FQErgo algorithm. Firstly, the all-quantum nature of the algorithm: unlike the previous methods [45, 46] for estimating ergotropy, FQErgo requires no classical optimizer. Secondly, FQErgo is robust against circuit errors since each feedback iteration readjusts the drives based on previous errors. This way, cumulative error growth is prevented. Thirdly, the circuit parameters can be read efficiently via an ancillary probe qubit using the interferometric method [47, 48, 49]. This way alleviates the need for resource-intensive quantum state tomography. Lastly, being a feedback-based method, FQErgo can be fully automated for convenient practical realizations. In this work, after numerically analyzing the FQErgo algorithm, we experimentally demonstrate it using two and three-qubit NMR registers.
@@ II.2 FQErgo: A feedback algorithm for ergotropy estimation
FQErgo is based on the FALQON algorithm proposed recently [42]. We assume multiple copies of individually accessible systems each prepared in the state $\rho_{0}$, which is not necessarily known. Alternatively, we assume a quantum system that can be repeatedly prepared into the same unknown state $\rho_{0}$. On introducing a drive Hamiltonian $\beta(t)H_{d}$, the system evolves as
$\frac{d}{dt}\rho(t)=i\left[H_{0}+\beta(t)H_{d}\,,\,\rho(t)\right],$ (4)
where $\beta(t)$ is the time-dependent drive parameter. Here and in the rest of the paper we have set $\hbar=1$. We seek to minimize the system energy $E(\rho(t))=\left\langle{H_{0}}\right\rangle_{\rho(t)}=\mathrm{Tr}(H_{0}\rho(t))$, which can be accomplished by designing control $\beta(t)$ such that
$\forall~{}t\geq 0,~{}\frac{d}{dt}\left\langle{H_{0}}\right\rangle_{\rho(t)}=\beta(t)\,\left\langle{C_{d}}\right\rangle_{\rho(t)}\leq 0,$ (5)
@@
Figure 1: (a) Illustrating FQErgo taking an initial state $\rho_{0}$ in $n$ iterations to $\rho_{n}\approx\rho^{p}$, its passive state. (b,c) Determining FQErgo drive amplitude $\beta$ with a probe qubit, for unitary Hermitian drive $\sigma_{\gamma}$ (b), and nonunitary Hermitian drive $A$ (c). For energy measurement, we set $A=H_{0}$.
@@ II.3 FQErgo using a probe qubit
Thus, we directly obtain the desired expectation value of the system as the signal $\left\langle{\sigma_{x}}\right\rangle_{\mathrm{probe}}$ in the probe qubit. For example, if the system Hamiltonian is the Pauli operator $\sigma_{z}$ as described later in the experimental section, the probe signal directly yields the expectation value $\left\langle{\sigma_{z}}\right\rangle_{\rho}$. For the more general case of extracting the expectation value $\left\langle{U}\right\rangle_{\rho}$ of a nonunitary Hermitian observable $A$, we construct a unitary $e^{-i\alpha A}$, where $\alpha$ is a scalar parameter such that $||\alpha A||\ll 1$. Now implementing the controlled gate as shown in Fig. 1 (c), we obtain the probe signal
@@ III.1 Single qubit system
Figure 2: (a,b) Simulated system-energy $E(\rho_{k})$ vs iteration number $k$ for a single qubit ($\omega_{0}=1)$ starting from 20 random initial states with ideal FQErgo (a) and FQErgo with random errors (b). Insets show the first three iterations. (c) Numerically estimated vs exact ergotropy values with ideal FQErgo (filled circles) and FQErgo with random errors (filled triangles).
Now we shall analyze the robustness of FQErgo against potential imperfections in practical implementations. Firstly, being a feedback-based method, FQErgo is very robust against the amplitude of the drives. The smaller the drive amplitude, it simply takes longer to reach the passive state. To study the robustness against more general errors, we introduce now an error rotation of 5 degrees about a random direction in every FQErgo iteration. The system-energy curves shown in Fig. 2 (b) do not show perfect monotonic decay, but neither exhibit serious build-up of errors. The corresponding ergotropy values shown by filled triangles in Fig. 2 (c) show somewhat underestimated values, which is expected since circuit errors can make the system settle in only a higher energy state by preventing it from reaching the passive state.
@@ III.2 Two qubit system
For two or more qubits, energy can be extracted using local unitaries on individual systems or global unitaries on the whole system. If we minimize the system energy using only local unitaries, we can get to the local passive state $\rho^{p_{L}}$, which determines the local ergotropy ${\cal E}^{L}=E(\rho_{0})-E(\rho^{p_{L}})$. If we minimize the system energy using global unitaries, along with or without local unitaries, we can get to the global passive state $\rho^{p_{G}}$, which determines the global ergotropy ${\cal E}^{G}=E(\rho_{0})-E(\rho^{p_{G}})\leq E(\rho^{p_{L}})$. The difference $\Delta={\cal E}^{G}-{\cal E}^{L}$ is called the ergotropy gap [50].
@@
Figure 3: (a,b) FQErgo circuits for local (a) and global (b) energy extraction in two-qubits. Here, local drives $X$, $Y$ suffice for reaching the local passive state $\rho^{(12)L}$ and estimating the local ergotropy ${\cal E}^{L}(\rho^{(12)}_{0})$ (a). In contrast, additional global drive such as $XY$ and $Z_{\delta}$ [45] gates are necessary for reaching the global passive state $\rho^{(12)G}$ and estimating the global ergotropy ${\cal E}^{G}(\rho^{(12)}_{0})$ (b). (c,d) Simulated system-energy vs iteration number for the two-qubit ($\omega_{0}=1$ and $J=1$) FQErgo from 20 random initial states without error (c) and with error (d). (e,f) The numerically estimated local (e) and global ergotropy (f) vs the exact values. (g) Ergotropy gap obtained from (e) and (f), vs the exact values. In (e-g), the filled circles are without random errors and filled triangles are with random errors.
@@
$U_{xy}=e^{-i\beta_{xy}H_{xy}\tau_{xy}}~{}\mbox{with}~{}H_{xy}=\sigma_{x}\otimes\sigma_{y}+\sigma_{y}\otimes\sigma_{x},$
strength $\beta_{xy}$, and duration $\tau_{xy}$. Finally, we find the second energy saturation over 60 iterations while reaching $\rho_{60}\approx\rho^{(12)p_{G}}$, the global passive state. Fig. 3 (c) shows the monotonic energy decrease for 20 random initial states undergoing FQErgo iteration. For each initial state, we observe two minimum-energy states, one corresponding to the local passive state and the other to the global passive state. Figs. 3 (e-f) plot the estimated local and global ergotropies against their exact values, and Fig. 3 (g) plots the estimated ergotropy gaps against exact values.
We now study FQErgo robustness by introducing a random nonlocal error unitary $U_{err}=e^{-iH_{err}\eta}$ generated by a random unit-norm error Hamiltonian $H_{err}$ and $\eta\equiv 2$ degrees in every FQErgo iteration. The system-energy curves shown in Fig. 3 (d) lose monotonicity in decay, but retain the overall trend. The error-affected local and global ergotropy values and the corresponding ergotropy gaps are shown by filled triangles in Fig. 3 (e-g). Likewise in one qubit case, the ergotropy values, particularly the global ergotropy values are underestimated, but the rms deviation of the ergotropy gap estimations from the ideal values remains below 0.07 indicating reasonable robustness of FQErgo against local/nonlocal errors.
@@ IV.1 Single-qubit ergotropy
We now use sodium fluorophosphate (dissolved in D2O; Fig. 4(a)) as the two-qubit register, wherein 19F is our probe qubit and 31P is the system qubit. All the experiments were carried out on a $500$ MHz Bruker NMR spectrometer at an ambient temperature of $300$ K. The rotating frame Hamiltonian of the system consisting of the internal part and the RF drive is $H^{\mathrm{FP}}=H^{\mathrm{FP}}_{\mathrm{int}}+H^{\mathrm{FP}}_{\mathrm{RF}}(t)$, where $H^{\mathrm{FP}}_{\mathrm{int}}=-\omega_{\mathrm{F}}I_{z}^{\mathrm{F}}-\omega_{\mathrm{P}}I_{z}^{\mathrm{P}}+2\pi J_{\mathrm{FP}}I_{z}^{\mathrm{F}}I_{z}^{\mathrm{P}}$ and $H^{\mathrm{FP}}_{\mathrm{RF}}(t)=\Omega_{\mathrm{F}}(t)I_{x}^{\mathrm{F}}+\Omega_{\mathrm{P}}(t)I_{x}^{\mathrm{P}}$, where $I_{x,y,z}$ are the spin operators. Here $(\omega_{\mathrm{F}}$, $\omega_{\mathrm{P}})$ and ($\Omega_{\mathrm{F}}(t)$, $\Omega_{\mathrm{P}}(t)$) respectively denote the adjustable frequency offsets and time-dependent RF amplitudes of 19F and 31P spins, and $J_{\mathrm{FP}}=868.0$ Hz is the scalar coupling constant.
@@
Figure 4: (a) Sodium fluorophosphate molecular structure with $J_{\mathrm{FP}}$ indicated. (b) The experimentally measured one-qubit energy (normalized) vs FQErgo iterations for three initial states $\rho_{\theta,\phi}=U_{\theta,\phi}|{00}\rangle\langle{00}|U^{\dagger}_{\theta,\phi}$ prepared by one of the RF rotations $U_{\theta,\phi}=e^{-i\theta(I_{x}\cos\phi+I_{y}\sin\phi)}\equiv(\theta,\phi)$ as indicated. (c) Ergotropy estimated from (b) vs simulated values. (d) Decaying fidelities of initial states (open symbols) and corresponding growth of passive states (filled symbols) for the same three cases as in (b).
@@
The pulse sequences for the NMR implementations of FQErgo are described in Appendix. Starting from the thermal state, we prepare a pseudopure state (PPS) before energizing the system qubit 31P into one of the three initial states $\rho_{0}$ as described in Fig. 4. For FQErgo, we choose the drives $I_{x}$ and $I_{y}$, and their strengths in the $k$th iteration are given by $\beta_{x(y)}=\left\langle{i[H_{0},I_{x(y)}]}\right\rangle_{\rho_{k-1}}$. Notice that all three different states reach their common passive state within 10 iterations in Fig. 4 (b). The corresponding ergotropy values show excellent agreement with the numerically simulated values in Fig. 4 (c), thus demonstrating the successful implementation of FQErgo.
@@
In this one qubit case, since we have now measured all three expectation values $\left\langle{\sigma_{x,y,z}}\right\rangle_{\rho_{k}}$, we can now reconstruct the density matrix and determine its state fidelities $F(\rho_{k})=\mathrm{Tr}(\rho_{0}\rho_{k})$ with the respective initial state and $F(\rho_{k})=\mathrm{Tr}(\rho_{0}\rho_{p})$ with the expected target passive state, at each iteration. The resulting fidelity profiles are shown in Fig. Fig. 4 (d). They indicate the gradual decay of the initial state fidelity and the corresponding buildup of the passive state fidelity. We clearly find high-fidelity passive states being prepared in each of the three cases.

### v2

@@ I Introduction
Quantum technologies are appealing as they can surpass their classical counterparts by using exclusive concepts like coherence and entanglement. Understanding their capabilities has progressed in recent years, thanks to the numerous experimental breakthroughs that have improved control over quantum states. Specifically, the investigation of the principles of quantum thermodynamics is made possible by quantum thermal machines, which include refrigerators and heat engines to regulate heat flow and work production [1, 2]. Further options for storing energy for later extraction include quantum batteries [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. The quantum batteries have the potential to outperform their classical counterparts in relevant metrics like charging speed [14, 15, 16, 17, 18, 19, 20], stored energy [21, 22, 20], and energy extraction [23, 24, 25, 26, 27, 28, 29].
@@
The main focus of this work is FQErgo, an algorithm we propose for preparing the passive state and thereby estimating ergotropy. FQErgo is inspired by (hence the name) the feedback-based algorithm for quantum optimization (FALQON) [44], which has been recently developed for combinatorial quantum optimization. This method is related to quantum Lyapunov control (QLC) and uses feedback conditioned on qubit measurements at each quantum circuit layer and parametrized quantum circuits to determine the values of the circuit parameters at subsequent layers [45, 46].
We note several salient features of the FQErgo algorithm. Firstly, the all-quantum nature of the algorithm: unlike the previous methods [47, 48] for estimating ergotropy, FQErgo requires no classical optimizer, and can be fully automated for convenient practical realizations. Secondly, FQErgo is robust against circuit errors since each feedback iteration readjusts the drives based on previous errors. This way, cumulative error growth is prevented. Thirdly, the circuit parameters can be read efficiently via an ancillary probe qubit using the interferometric method [49, 50, 51], which alleviates the need for resource-intensive quantum state tomography. In this work, after numerically analyzing the FQErgo algorithm, we experimentally demonstrate it using two and three-qubit NMR registers.
@@ II.2 FQErgo: A feedback algorithm for ergotropy estimation
Figure 1: (a) Illustrating $n$ iterations of FQErgo applied on an initial state $\rho_{0}$ to reach $\rho_{n}\approx\rho^{p}$, its passive state. (b,c) Determining FQErgo drive amplitude $\beta$ with a probe qubit, for (b) unitary Hermitian drive $\sigma_{\gamma}$, and (c) nonunitary Hermitian drive $A$. For energy measurement, we set $A=H_{0}$.
FQErgo is based on the FALQON algorithm proposed recently [44]. We assume multiple copies of individually accessible systems each prepared in the state $\rho_{0}$, which is not necessarily known. Alternatively, we can assume an oracle that transforms a uniquely initialized quantum system into a definite unknown state $\rho_{0}$. On introducing a drive Hamiltonian $\beta(t)H_{d}$, the system evolves as
$i\frac{d}{dt}\rho(t)=\left[H_{0}+\beta(t)H_{d}\,,\,\rho(t)\right],$ (4)
where $\beta(t)$ is the time-dependent drive parameter. Here and in the rest of the paper we have set $\hbar=1$. We seek to minimize the system energy $E(\rho(t))=\left\langle{H_{0}}\right\rangle_{\rho(t)}=\mathrm{Tr}(H_{0}\rho(t))$, which can be accomplished by designing control $\beta(t)$ satisfying the Lyapunov condition
$\frac{d}{dt}\left\langle{H_{0}}\right\rangle_{\rho(t)}=\beta(t)\,\left\langle{C_{d}}\right\rangle_{\rho(t)}\leq 0,~{}~{}\forall~{}t\geq 0$ (5)
@@ II.3 FQErgo using a probe qubit
Thus, we directly obtain the desired expectation value of the system as the signal $\left\langle{\sigma_{x}}\right\rangle_{\mathrm{probe}}$ in the probe qubit. For example, if the system Hamiltonian is the Pauli operator $\sigma_{z}$ as described later in the experimental section, the probe signal directly yields the expectation value $\left\langle{\sigma_{z}}\right\rangle_{\rho}$. For the more general case of extracting the expectation value $\left\langle{U}\right\rangle_{\rho}$ of a nonunitary Hermitian observable $A$, we construct a unitary $e^{-i\alpha A}$, where $\alpha$ is a real parameter such that $||\alpha A||\ll 1$. Now implementing the controlled gate as shown in Fig. 1 (c), we obtain the probe signal [50]
@@
Figure 2: (a,b) Simulated system-energy $E(\rho_{k})$ vs iteration number $k$ for a single qubit ($\omega_{0}=1)$ starting from 20 random initial states applied with (a) ideal FQErgo and (b) FQErgo having random errors. Insets show the first three iterations. (c) Numerically estimated vs exact ergotropy values with ideal FQErgo (filled circles) and FQErgo with random errors (filled triangles).
@@ III.1 Single qubit system
Now we numerically analyze the robustness of FQErgo against potential imperfections in practical implementations. Firstly, being a feedback-based method, FQErgo is naturally robust against the drive amplitude. If the drive amplitude is lesser than the nominal, it still reaches the passive state but simply takes more iterations. To study the effect of more general errors, we introduce an error rotation of 5 degrees about a random direction in every FQErgo iteration. The system-energy curves in Fig. 2 (b) no longer exhibit a perfect monotonic decay, but neither show serious build-up of errors. The corresponding ergotropy values shown by filled triangles in Fig. 2 (c) show somewhat underestimated values, which is expected since circuit errors can only make the system settle in a higher energy state by preventing it from reaching the passive state.
Figure 3: (a,b) FQErgo circuits for local (a) and global (b) energy extraction in two-qubits. Here, local drives $X$, $Y$ suffice for reaching the local passive state $\rho^{(12)L}$ and estimating the local ergotropy ${\cal E}^{L}(\rho^{(12)}_{0})$ (a). In contrast, additional global drive such as $XY$ and $Z_{\delta}$ [47] gates are necessary for reaching the global passive state $\rho^{(12)G}$ and estimating the global ergotropy ${\cal E}^{G}(\rho^{(12)}_{0})$ (b). (c,d) Simulated system-energy vs iteration number for the two-qubit ($\omega_{0}=1$ and $J=0.01$) FQErgo from 20 random initial states without error (c) and with error (d). (e,f) The numerically estimated local (e) and global ergotropy (f) vs the exact values. (g) Ergotropy gap obtained from (e) and (f), vs the exact values. In (e-g), the filled circles are without random errors and filled triangles are with random errors.
@@ III.2 Two qubit system
There is a fundamental difference between work extractions from one-qubit and multi-qubit systems. For two or more qubits, energy can be extracted using local unitaries on individual systems or global unitaries on the whole system. If we minimize the system energy using only local unitaries, we can get to the local passive state $\rho^{p_{L}}$, which determines the local ergotropy ${\cal E}^{L}=E(\rho_{0})-E(\rho^{p_{L}})$. If the minimization controls involve global unitaries, in addition to local, we can get to the global passive state $\rho^{p_{G}}$, which determines the global ergotropy ${\cal E}^{G}=E(\rho_{0})-E(\rho^{p_{G}})\leq E(\rho^{p_{L}})$. The difference $\Delta={\cal E}^{G}-{\cal E}^{L}$ is called the ergotropy gap [52].
@@
$U_{xy}=e^{-i\beta^{xy}H_{xy}\tau_{xy}}~{}\mbox{with}~{}H_{xy}=\sigma_{x}\otimes\sigma_{y}+\sigma_{y}\otimes\sigma_{x},$
strength $\beta^{xy}$, and duration $\tau_{xy}$. Finally, we find the second energy saturation over 60 iterations while reaching $\rho_{60}\approx\rho^{(12)p_{G}}$, the global passive state. Fig. 3 (c) shows the monotonic energy decrease for 20 random initial states undergoing FQErgo iteration. For each initial state, we observe two minimum-energy states, one corresponding to the local passive state and the other to the global passive state. Figs. 3 (e-f) plot the estimated local and global ergotropies against their exact values, and Fig. 3 (g) plots the estimated ergotropy gaps against exact values. Indeed, one can apply global energy extraction sequence Fig. 3 (b) directly from initial point and reach passive state much faster but here our focus is to estimate the ergotropy gap $\Delta$.
We again study FQErgo robustness by introducing a random nonlocal error unitary $U_{err}=e^{-iH_{err}\eta}$ generated by a random unit-norm error Hamiltonian $H_{err}$ and $\eta\equiv 2$ degrees in every FQErgo iteration. The system-energy curves shown in Fig. 3 (d) lose monotonicity in decay, but retain the overall trend. The error-affected local and global ergotropy values and the corresponding ergotropy gaps are shown by filled triangles in Fig. 3 (e-g). Likewise in one qubit case, the ergotropy values, particularly the global ergotropy values are underestimated, but the rms deviation of the ergotropy gap estimations from the ideal values remains below 0.07 indicating reasonable robustness of FQErgo against unitary errors.
@@ IV.1 Single-qubit ergotropy
We now use sodium fluorophosphate (dissolved in D2O; Fig. 4(a)) as the two-qubit register, wherein 19F nuclear spin is our probe qubit and 31P nuclear spin is the system qubit. All the experiments were carried out on a $500$ MHz Bruker NMR spectrometer at an ambient temperature of $300$ K. The rotating frame Hamiltonian of the system consisting of the internal part and the RF drive is $H^{\mathrm{FP}}=H^{\mathrm{FP}}_{\mathrm{int}}+H^{\mathrm{FP}}_{\mathrm{RF}}(t)$, where $H^{\mathrm{FP}}_{\mathrm{int}}=-\omega_{\mathrm{F}}I_{z}^{\mathrm{F}}-\omega_{\mathrm{P}}I_{z}^{\mathrm{P}}+2\pi J_{\mathrm{FP}}I_{z}^{\mathrm{F}}I_{z}^{\mathrm{P}}$ and $H^{\mathrm{FP}}_{\mathrm{RF}}(t)=\Omega_{\mathrm{F}}(t)I_{x}^{\mathrm{F}}+\Omega_{\mathrm{P}}(t)I_{x}^{\mathrm{P}}$, where $I_{x,y,z}$ are the spin operators. Here $(\omega_{\mathrm{F}}$, $\omega_{\mathrm{P}})$ and ($\Omega_{\mathrm{F}}(t)$, $\Omega_{\mathrm{P}}(t)$) respectively denote the adjustable frequency offsets and time-dependent RF amplitudes of 19F and 31P spins, and $J_{\mathrm{FP}}=868.0$ Hz is the scalar coupling constant.
@@
Figure 4: (a) Sodium fluorophosphate molecular structure with $J_{\mathrm{FP}}$ indicated. (b) The experimentally measured one-qubit energy (normalized) vs FQErgo iterations for three initial states $\rho_{\theta,\phi}=|{0}\rangle\langle{0}|\otimes U_{\theta,\phi}|{0}\rangle\langle{0}|U^{\dagger}_{\theta,\phi}$ prepared by one of the RF rotations $U_{\theta,\phi}=e^{-i\theta(I_{x}\cos\phi+I_{y}\sin\phi)}\equiv(\theta,\phi)$ as indicated. (c) Ergotropy estimated from (b) vs simulated values. (d) Decaying fidelities of initial states ($F_{0}(\rho_{k})$, open symbols) and corresponding growth of passive states ($F_{p}(\rho_{k})$, filled symbols) for the same three cases as in (b).
@@
The pulse sequences for the NMR implementations of FQErgo are described in Appendix. Starting from the thermal state, we prepare a pseudopure state (PPS) before energizing the system qubit 31P into one of the three initial states $\rho_{0}$ as described in Fig. 4. For FQErgo, we randomly choose the initial drives $\beta^{x}_{0}I_{x}$ and $\beta^{y}_{0}I_{y}$, and their strengths in the $k$th iteration are obtained by measuring the ancilla signal $\left\langle{i[H_{0},I_{x(y)}]}\right\rangle_{\rho_{k-1}}=\beta^{x(y)}_{k}$. Notice that all three different states reach their common passive state within 10 iterations in Fig. 4 (b). The corresponding ergotropy values show excellent agreement with the numerically simulated values in Fig. 4 (c), thus demonstrating the successful implementation of FQErgo.
@@
In this one qubit case, since we have now measured all three expectation values $\left\langle{\sigma_{x,y,z}}\right\rangle_{\rho_{k}}$, we can now reconstruct the density matrix and determine its state fidelities $F_{0}(\rho_{k})=\mathrm{Tr}(\rho_{0}\rho_{k})$ with the respective initial state, and $F_{p}(\rho_{k})=\mathrm{Tr}(\rho_{p}\rho_{k})$ with the expected target passive state, at each iteration. The resulting fidelity profiles are shown in Fig. Fig. 4 (d). They indicate the gradual decay of the initial state fidelity and the corresponding buildup of the passive state fidelity. We clearly find high-fidelity passive states being prepared in each of the three cases.